OpenStudy (anonymous):

if the point (2/3, y) is on the unit circle, what is y? A.) 2/3 B.) 1/3 C.) √5/3 D.) 4/9

OpenStudy (anonymous):

Do you know what algebraic equation describes the unit circle?

OpenStudy (anonymous):

no i don't

OpenStudy (welshfella):



OpenStudy (anonymous):

Ok, the unit circle is described by the equation\[x^2 + y^2 = 1\]

OpenStudy (welshfella):

its that man Pythagoras again..

OpenStudy (anonymous):

So, if x = 2/3, then we can compute\[y^2 = 1 - x^2 \implies y^2 = 1 - \left(\frac{2}{3}\right)^2 \]

OpenStudy (anonymous):

Yes, Pythagoras, he gets around.

OpenStudy (anonymous):

haha okay I got it now ^_^ thank you guys so much!